A rectangular room is 1.8 times as long as it is wide, and its perimeter is 30 meters. Find the dimension of the room.

The length is: meters and the width is: meters.

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND YOU CAN DO YOUR PROBLEM YOUR SELF!!!!
perimeter of rectangle is 64 centimeters length is 4 centimeters less than twice the width what is lenght and width? thanks
let the width be x..twice width=2x...4 cm. less than that =2x-4
perimeter =2(length+width)=64
2(x+2x-4)=64
2(3x-4)=64
6x-8=64
6x=64+8=72
x=72/6=12
so width is 12 cm and length is 2*12-4=20 cm.